OpenStudy (anonymous):

Express the volume of the rectangular prism in simplest radical form. l: 5√2 + 2√3, w: √15 - √12 and h: 5√2 - 2√3?

OpenStudy (mertsj):

\[(5\sqrt{2}+2\sqrt{3})(\sqrt{15}-2\sqrt{3})(5\sqrt{2}-2\sqrt{3})\] Let's rearrange for convenience.

OpenStudy (mertsj):

\[(5\sqrt{2}+2\sqrt{3})(5\sqrt{2}-2\sqrt{3})(\sqrt{15}-2\sqrt{3})\] \[(50-12)(\sqrt{15}-2\sqrt{3})\] \[38(\sqrt{15}-2\sqrt{3})=38\sqrt{15}-76\sqrt{3}\]

OpenStudy (anonymous):

the answer has to be: 38√15 - 38√2

OpenStudy (mertsj):

Well then you have posted something wrong because the answer I gave you is the correct one for the problem you posted.

OpenStudy (anonymous):



OpenStudy (mertsj):